if perimeter of rectangle is 2oo than what is the area of rectangle? solve
Answer by Alan3354(69443)

if perimeter of rectangle is 2oo than what is the area of rectangle? solve
--------------
P = SL + 2W = 200
L + W = 100
L*W = Area